VIBRATIONAL DYNAMICS OF TRIFLUOROACETIC ACID AND FORMIC ACID IN GAS AND DILUTE SOLUTION: CRACKING OPEN GAS PHASE ACID DIMERS

\underline{PAM L. CRUM}, BRIAN C. DIAN, HYUN S. YOO, CHARLOTTE E. HINKLE AND BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904..

Ultrafast time-domain transient absorption spectroscopy is used to study the vibrational dynamics of the cyclic hydrogen bonded dimers of trifluoroacetic acid and formic acid in both the gas- and solution-phase. Deposition of energy into the broad dimer bands of these acid dimers in the gas-phase provides the means to break open the dimer, as evidenced by the growth of a ` ` free O-H´ ´ absorbance at ˜ 3580cm^{−1}on the time scales of energy flow out of the initially excited dimer region. In trifluoroacetic acid there is no change in the rate of ` ` free O-H´ ´ growth with varying pump frequency between 2596cm^{−1}and 3131cm^{−1}. Spectrally resolved pump-probe experiments are performed to reveal the evolution of a broad range of the spectra in time, which facilitates interpretation of the dynamics behind the time scales provided by single frequency measurements. When the acid dimers are solvated (0.05M CCl_{4} solution), energy pumped into the broad dimer band can be cooled out by the solvent, making it unavailable to the process of dimer opening.